# S2malloc: Statistically Secure Allocator for Use-After-Free Protection And More

### v1

@@ 6.3. Performance on real world programs
Throughput Memory (KB) p50 (msec) Throughput Memory (KB) p50 (msec)
@@
Table 7. Memory consumption, throughput, and delays for servers
@@
Throughput Memory (KB) Throughput Memory (KB) Throughput Memory (KB)
@@
Table 8. Memory Consumption and Throughput on databases
6.4. Performance with multi-threading
@@
6.5. Influence with different parameters
@@ 8. Conclusion
9. Acknowledgment
This work is funded in part by NSERC (RGPIN-2022-03325), the David R. Cheriton endowment, and gifts from Facebook, Amazon, Sui Foundation, and Intel Labs.
Appendix A List of failed tests and corresponding exceptions
@@
Appendix B Run time and memory overheads of each PARSEC and SPEC test
Setup Run-time overhead (normalized with default allocator) Memory overhead (normalized with default allocator)
@@ Appendix B Run time and memory overheads of each PARSEC and SPEC test
Table 9. Average and standard deviation of run time and memory overhead on PARSEC and SPEC benchmarks (x86 and AARCH).
Appendix C Run time and memory overheads of each mimalloc-bench test
@@
Table 10. Normalized runtime overheads of mimalloc-bench
@@
Table 11. Normalized memory overheads of mimalloc-bench
Appendix D On The Formal Modeling of Probabilistic Use-After-Free Detection
@@
D.1. Success rate of attack and defense per single attempt
@@
(1) $A=\dfrac{1}{r}\vartriangleright\text{\footnotesize{the correct block}}\vartriangleleft\times\dfrac{1}{1+\lfloor\dfrac{b-s}{16}\rfloor}\vartriangleright\text{\footnotesize{the correct in-block RIO value}}\vartriangleleft=\dfrac{1}{r(1+\lfloor\dfrac{b-s}{16}\rfloor)}$
@@
(2) $D=\dfrac{2\sum_{\_}{i=0}^{c-2}(l+i)+\sum_{\_}{\_={c-1}}^{b-(l+c-1)}(l+c-1)\vartriangleright\text{\footnotesize{number of overlaps}}\vartriangleleft}{(b-l+1)(b-c+1)\vartriangleright\text{\footnotesize{number ofways to place $l$-byte and $c$-byte}}\vartriangleleft}=\dfrac{b(l+c-1)-(l-1)^{2}-(c-1)^{2}-cl+1}{(b-l+1)(b-c+1)}$
@@
D.2. Strategy S1: repetitive UAF-writes to the same address
@@
(4) _$\mathbb{P}_{\_}{e}^{i+1}=\mathbb{P}_{\_}{e}^{i}\vartriangleright\text{\footnotesize{reaches round $i$}}\vartriangleleft\times(1-\dfrac{2d+1}{r}\times D)\vartriangleright\text{\footnotesize{undetected}}\vartriangleleft\times(1-A)\vartriangleright\text{\footnotesize{unsuccessful attack attempt}}\vartriangleleft$
@@
(5) _$\mathbb{P}_{\_}{attack}^{K}=A+\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times A)\quad\quad::\quad\quad\mathbb{P}_{\_}{detect}^{K}=\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times\dfrac{2d+1}{r}\times D)$
D.3. Strategy S2: UAF-writes through fresh dangling pointers
Unlike § D.2, the attacker does not hold a dangling pointer indefinitely, instead, the attacker obtains a fresh dangling pointer ( 2) if a prior UAF-write attempt is not successful. After obtaining a dangling pointer, if the attacker notices a victim object is allocated ( 4), an $l$-byte UAF-write through the fresh dangling pointer is conducted ( 5). This essentially means that every UAF-write is likely on a different address.
@@ D.3. Strategy S2: UAF-writes through fresh dangling pointers
We use round $i$ to represent the $i$-th round of the game. In each round, the attacker makes the first move and S2malloc follows. We denote _$\mathbb{P}_{\_}{e}^{i}$ to represent the probability that the program execution ever reaches S2malloc’s turn in round $i$ (to be consistent with the notation in § D.2). By definition, _$\mathbb{P}_{\_}{e}^{1}=1-A$, i.e., the attacker’s first UAF-write is not successful in achieving the goal ( 1).
@@
(6) _$\mathbb{Q}_{\_}i=(\dfrac{r-1}{r}\vartriangleright\text{\footnotesize{ball not selected}}\vartriangleleft+\dfrac{1}{r}\times(1-D)\vartriangleright\text{\footnotesize{ball selected but not colored}}\vartriangleleft)^{i}=(\dfrac{r-D}{r})^{i}$
@@
(7) _$\mathbb{P}_{\_}{d}^{i}=1-\dfrac{Q_{\_}ir-2d}{r-2d}\cdot\prod_{\_}{i=0}^{2d}(\dfrac{Q_{\_}ir-i}{r-i})$
@@
(8) _$\mathbb{P}_{\_}{e}^{i+1}=\mathbb{P}_{\_}{e}^{i}\vartriangleright\text{\footnotesize{reaches round $i$}}\vartriangleleft\times(1-\mathbb{P}_{\_}{d}^{i})\vartriangleright\text{\footnotesize{undetected}}\vartriangleleft\times(1-A)\vartriangleright\text{\footnotesize{unsuccessful attack attempt}}\vartriangleleft$
@@
(9) _$\mathbb{P}_{\_}{attack}^{K}=A+\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times A)\quad\quad::\quad\quad\mathbb{P}_{\_}{detect}^{K}=\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times\mathbb{P}_{\_}{d}^{i})$
D.4. Strategy S1-spray: repetitive UAF-writes to the same address with spraying
This strategy operates similarly to the strategy in § D.2: the attacker first obtains a dangling pointer ( 2) and holds the pointer for arbitrarily long. However, generalized from § D.2 in which the attacker conducts a UAF-write after one allocation of a victim object ( 4), the attacker waits until there are $m$ victim objects newly allocated and alive (i.e., not freed yet) and then issues the UAF-write. Effectively, the attacker is trying to diligently spray the heap by victim objects to increase its chance of success. The UAF-write is still an $l$-byte memory write at the same offset through the same dangling pointe ( 5). This is essentially repetitive UAF-writes to the same address, similar to § D.2.
Consistent with the analysis in § D.2, we still use round $i$ to represent the $i$-th allocation of a victim object after the attacker obtains the dangling pointer and conducts the UAF-write. Effectively, after round $i$, S2malloc should have checked FBCs $i$ times to catch the UAF attempt.
@@ D.4. Strategy S1-spray: repetitive UAF-writes to the same address with spraying
And yet, consistent with § D.2, this UAF-write can corrupt one FBC at most. Hence, suppose the execution has reached round $i$, the probability that the repetitive UAF-writes is detected at this particular round is still Equation 3. With attacker’s success rate improved, the inductive definition for _$\mathbb{P}_{\_}{e}^{i}$ in this strategy will be:
(10) _$\mathbb{P}_{\_}{e}^{i+1}=\mathbb{P}_{\_}{e}^{i}\vartriangleright\text{\footnotesize{reaches round $i$}}\vartriangleleft\times(1-\dfrac{2d+1}{r}\times D)\vartriangleright\text{\footnotesize{undetected}}\vartriangleleft\times(1-mA)\vartriangleright\text{\footnotesize{unsuccessful attack attempt}}\vartriangleleft$
@@
(11) _$\mathbb{P}_{\_}{attack}^{K}=mA+\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times mA)\quad\quad::\quad\quad\mathbb{P}_{\_}{detect}^{K}=\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times\dfrac{2d+1}{r}\times D)$
D.5. Strategy S2-spray: UAF-writes through fresh dangling pointers with spraying
In this strategy, the attacker still sprays the heap such that overriding the sensitive field in any the $m$ victim objects achieves the goal (like the strategy in § D.4). Similar to § D.3, the attacker does not hold a dangling pointer indefinitely, instead, the attacker obtains a fresh dangling pointer ( 2) if a prior UAF-write attempt is not successful and use it to launch a UAF-write attack in the current round. In each round, the attacker makes the first move and S2malloc follows. We denote _$\mathbb{P}_{\_}{e}^{i}$ to represent the probability that the program execution ever reaches S2malloc’s turn in round $i$ (to be consistent with the notation in § D.2§ D.3, and § D.4). By definition, _$\mathbb{P}_{\_}{e}^{1}=1-mA$, i.e., the attacker’s first UAF-write is not successful in achieving the goal ( 1) even after spraying $m$ victim objects.
Similar to § D.3, we use _$\mathbb{Q}_{\_}i$ to denote the probability that an arbitrary ball in the box is not colored (i.e., a block with its FBCs integral) after $i$ rounds. _$\mathbb{Q}_{\_}i$ and _$\mathbb{P}_{\_}{d}^{i}$ can be computed using the same formula as in § D.3
@@ D.5. Strategy S2-spray: UAF-writes through fresh dangling pointers with spraying
(12) _$\mathbb{P}_{\_}{e}^{i+1}=\mathbb{P}_{\_}{e}^{i}\vartriangleright\text{\footnotesize{reaches round $i$}}\vartriangleleft\times(1-\mathbb{P}_{\_}{d}^{i})\vartriangleright\text{\footnotesize{undetected}}\vartriangleleft\times(1-mA)\vartriangleright\text{\footnotesize{unsuccessful attack attempt}}\vartriangleleft$
@@
(13) _$\mathbb{P}_{\_}{attack}^{K}=mA+\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times mA)\quad\quad::\quad\quad\mathbb{P}_{\_}{detect}^{K}=\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times\mathbb{P}_{\_}{d}^{i})$

### v2

@@ 6.3 Performance on real world programs
Throughput Memory p50 Throughput Memory p50
@@
Table 7: Memory consumption (KB), throughput (requests/second), and delays (msec) for servers
@@
Throughput Memory Throughput Memory Throughput Memory
@@
Table 8: Memory Consumption (KB) and throughput (requests/second) on databases
6.4 Performance with multi-threading
@@
6.5 Influence with different parameters
@@ 8 Conclusion
9 Acknowledgment
This work is funded in part by NSERC (RGPIN-2022-03325), the David R. Cheriton endowment, and gifts from Facebook, Amazon, Sui Foundation, and Intel Labs. We are grateful for the insightful feedback and guidance from our shepherd, Dr. R. Sekar which helped us to improve the paper significantly.
Appendix 0.A List of failed tests and corresponding exceptions
@@
Appendix 0.B Run time and memory overheads of each PARSEC and SPEC test
Setup Run-time overhead Memory overhead
@@ Appendix 0.B Run time and memory overheads of each PARSEC and SPEC test
Table 9: Average and standard deviation of run time and memory overhead on PARSEC and SPEC benchmarks (x86 and AARCH).
Appendix 0.C Run time and memory overheads of each mimalloc-bench test
@@
Table 10: Normalized runtime overheads of mimalloc-bench
@@
Table 11: Normalized memory overheads of mimalloc-bench
Appendix 0.D Adapted code snippets to illustrate CVE-2015-6835 and its exploits
1HashTable *SESSION_VARS;
2
3PS_SERIALIZER_DECODE_FUNC(char* p, char *endptr) {
4  char *cursor = p;
5  hash_table_t *ht = INIT_HASHTABLE();
6  while (cursor < endptr) {
7    // for each item in the serialized stream
8    zval *name = PARSE_NEXT_ZVAL(&cursor);
9
10    zval *parsed = php_var_unserialize(&cursor, ht, endptr);
11    zval *stored = emalloc(sizeof(zval));
12    COPY_ZVAL(parsed, stored);
13    SESSION_VARS[name->value.str.val] = stored; // refcount = 1
14
15    zval_dtor(name);
16    zval_dtor(parsed);
17  }
18  free(ht);
19}
20zval* php_var_unserialize(char** pp, hash_table_t* ht, char* endptr) {
21  char *cursor = *pp;
22  zval *rval = emalloc(sizeof(zval));
23  INIT_ZVAL(rval); // refcount = 1
24
25  while (cursor < endptr) {
26    // for each element in the serialized item
27    zval *tmp;
28    if (*cursor == R) {
29        // reference to a previously parsed zval
30        cursor++;
31        tmp = PARSE_NEXT_ZVAL(&cursor);
32        tmp = ht[HASH_ZVAL(tmp)];
33    } else {
34        // parsing a new zval from the cursor
35        tmp = PARSE_NEXT_ZVAL(&cursor);
36    }
37    rval->value.ht[HASH_ZVAL(tmp)] = tmp;
38    zval_dtor(tmp);
39  }
40
41  *pp = cur;
42  ht[HASH_ZVAL(rval)] = rval;
43  return rval;
44}
45void zval_dtor(zval *p) {
46    p->refcount--;
47    if (p->refcount == 0) { free(p); }
48}
49
50/* ---------------------------------------------------------------- */
51
52int zend_echo_handler(char *name) {
53  zval *obj = SESSION_VARS[name];
54  obj->refcount++;
55  zend_print_variable(obj); // UAF read
56}
Figure 7: Adapted code snippets to illustrate CVE-2015-6835 and its exploits.
Appendix 0.E On The Formal Modeling of Probabilistic Use-After-Free Detection
@@
0.E.1 Success rate of attack and defense per single attempt
@@
$A=\dfrac{1}{r}\vartriangleright\text{\small{the correct block}}\vartriangleleft\times\dfrac{1}{1+\lfloor\dfrac{b-s}{16}\rfloor}\vartriangleright\text{\small{the correct in-block RIO value}}\vartriangleleft=\dfrac{1}{r(1+\lfloor\dfrac{b-s}{16}\rfloor)}$ (1)
@@
$D$ $=\dfrac{2\sum_{\_}{i=0}^{c-2}(l+i)+\sum_{\_}{\_={c-1}}^{b-(l+c-1)}(l+c-1)\vartriangleright\text{\small{number of overlaps}}\vartriangleleft}{(b-l+1)(b-c+1)\vartriangleright\text{\small{number of ways to place $l$-byte and$c$-byte}}\vartriangleleft}$ (2)
$=\dfrac{b(l+c-1)-(l-1)^{2}-(c-1)^{2}-cl+1}{(b-l+1)(b-c+1)}$
@@
0.E.2 Strategy S1: repetitive UAF-writes to the same address
@@
_$\mathbb{P}_{\_}{e}^{i+1}=\mathbb{P}_{\_}{e}^{i}$ reaches round $\vartriangleright\text{\small{reaches round $i$}}\vartriangleleft\times(1-\dfrac{2d+1}{r}\times D)\vartriangleright\text{\small{undetected}}\vartriangleleft$ (4)
[…]
@@
_$\mathbb{P}_{\_}{attack}^{K}=A+\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times A)\quad\quad::\quad\quad\mathbb{P}_{\_}{detect}^{K}=\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times\dfrac{2d+1}{r}\times D)$ (5)
0.E.3 Strategy S2: UAF-writes through fresh dangling pointers
Unlike § 0.E.2, the attacker does not hold a dangling pointer indefinitely, instead, the attacker obtains a fresh dangling pointer ( 2) if a prior UAF-write attempt is not successful. After obtaining a dangling pointer, if the attacker notices a victim object is allocated ( 4), an $l$-byte UAF-write through the fresh dangling pointer is conducted ( 5). This essentially means that every UAF-write is likely on a different address.
@@ 0.E.3 Strategy S2: UAF-writes through fresh dangling pointers
We use round $i$ to represent the $i$-th round of the game. In each round, the attacker makes the first move and S2malloc follows. We denote _$\mathbb{P}_{\_}{e}^{i}$ to represent the probability that the program execution ever reaches S2malloc’s turn in round $i$ (to be consistent with the notation in § 0.E.2). By definition, _$\mathbb{P}_{\_}{e}^{1}=1-A$, i.e., the attacker’s first UAF-write is not successful in achieving the goal ( 1).
@@
_$\mathbb{Q}_{\_}i=(\dfrac{r-1}{r}\vartriangleright\text{\small{ball not selected}}\vartriangleleft+\dfrac{1}{r}\times(1-D)\vartriangleright\text{\small{ball selected but not colored}}\vartriangleleft)^{i}=(\dfrac{r-D}{r})^{i}$ (6)
@@
_$\mathbb{P}_{\_}{d}^{i}=1-\dfrac{Q_{\_}ir-2d}{r-2d}\cdot\prod_{\_}{i=0}^{2d}(\dfrac{Q_{\_}ir-i}{r-i})$ (7)
@@
_$\mathbb{P}_{\_}{e}^{i+1}=\mathbb{P}_{\_}{e}^{i}$ $\vartriangleright\text{\small{reaches round $i$}}\vartriangleleft\times(1-\mathbb{P}_{\_}{d}^{i})\vartriangleright\text{\small{undetected}}\vartriangleleft\times$ (8)
[…]
@@
_$\mathbb{P}_{\_}{attack}^{K}=A+\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times A)\quad\quad::\quad\quad\mathbb{P}_{\_}{detect}^{K}=\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times\mathbb{P}_{\_}{d}^{i})$ (9)
0.E.4 Strategy S1-spray: repetitive UAF-writes to the same address with spraying
This strategy operates similarly to the strategy in § 0.E.2: the attacker first obtains a dangling pointer ( 2) and holds the pointer for arbitrarily long. However, generalized from § 0.E.2 in which the attacker conducts a UAF-write after one allocation of a victim object ( 4), the attacker waits until there are $m$ victim objects newly allocated and alive (i.e., not freed yet) and then issues the UAF-write. Effectively, the attacker is trying to diligently spray the heap by victim objects to increase its chance of success. The UAF-write is still an $l$-byte memory write at the same offset through the same dangling pointe ( 5). This is essentially repetitive UAF-writes to the same address, similar to § 0.E.2.
Consistent with the analysis in § 0.E.2, we still use round $i$ to represent the $i$-th allocation of a victim object after the attacker obtains the dangling pointer and conducts the UAF-write. Effectively, after round $i$, S2malloc should have checked FBCs $i$ times to catch the UAF attempt.
@@ 0.E.4 Strategy S1-spray: repetitive UAF-writes to the same address with spraying
And yet, consistent with § 0.E.2, this UAF-write can corrupt one FBC at most. Hence, suppose the execution has reached round $i$, the probability that the repetitive UAF-writes is detected at this particular round is still Equation 3. With attacker’s success rate improved, the inductive definition for _$\mathbb{P}_{\_}{e}^{i}$ in this strategy will be:
_$\mathbb{P}_{\_}{e}^{i+1}=\mathbb{P}_{\_}{e}^{i}$ $\vartriangleright\text{\small{reaches round $i$}}\vartriangleleft\times(1-\dfrac{2d+1}{r}\times D)\vartriangleright\text{\small{undetected}}\vartriangleleft\times$ (10)
[…]
@@
_$\mathbb{P}_{\_}{attack}^{K}=mA+\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times mA)\quad\quad::\quad\quad\mathbb{P}_{\_}{detect}^{K}=\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times\dfrac{2d+1}{r}\times D)$ (11)
0.E.5 Strategy S2-spray: UAF-writes through fresh dangling pointers with spraying
In this strategy, the attacker still sprays the heap such that overriding the sensitive field in any the $m$ victim objects achieves the goal (like the strategy in § 0.E.4). Similar to § 0.E.3, the attacker does not hold a dangling pointer indefinitely, instead, the attacker obtains a fresh dangling pointer ( 2) if a prior UAF-write attempt is not successful and use it to launch a UAF-write attack in the current round. In each round, the attacker makes the first move and S2malloc follows. We denote _$\mathbb{P}_{\_}{e}^{i}$ to represent the probability that the program execution ever reaches S2malloc’s turn in round $i$ (to be consistent with the notation in § 0.E.2§ 0.E.3, and § 0.E.4). By definition, _$\mathbb{P}_{\_}{e}^{1}=1-mA$, i.e., the attacker’s first UAF-write is not successful in achieving the goal ( 1) even after spraying $m$ victim objects.
Similar to § 0.E.3, we use _$\mathbb{Q}_{\_}i$ to denote the probability that an arbitrary ball in the box is not colored (i.e., a block with its FBCs integral) after $i$ rounds. _$\mathbb{Q}_{\_}i$ and _$\mathbb{P}_{\_}{d}^{i}$ can be computed using the same formula as in § 0.E.3
@@ 0.E.5 Strategy S2-spray: UAF-writes through fresh dangling pointers with spraying
_$\mathbb{P}_{\_}{e}^{i+1}=\mathbb{P}_{\_}{e}^{i}$ $\vartriangleright\text{\small{reaches round $i$}}\vartriangleleft\times(1-\mathbb{P}_{\_}{d}^{i})\vartriangleright\text{\small{undetected}}\vartriangleleft\times$ (12)
[…]
@@
_$\mathbb{P}_{\_}{attack}^{K}=mA+\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times mA)\quad\quad::\quad\quad\mathbb{P}_{\_}{detect}^{K}=\sum_{\_}{i=1}^{K}(\mathbb{P}_{\_}{e}^{i}\times\mathbb{P}_{\_}{d}^{i})$ (13)